1/1

Lord George Gordon Byron ,  Lord George Gordon Byron ,
The harp the monarch minstrel swept  O Harfe, die des Gottgeliebten Hand

tradukita de Franz Theremin

The harp the monarch minstrel swept, The King of men, the loved of Heaven, Which Music hallow’d while she wept O’er tones her heart of hearts had given, Redoubled be her tears, its chords are riven! It soften’d men of iron mould, It gave them virtues not their own; No ear so dull, no soul so cold, That felt not, fired not to the tone, Till David’s lyre grew mightier than his throne!

O Harfe, die des Gottgeliebten Hand, Des königlichen Sängers, hat geschlagen, Die, was die Andacht Innigstes empfand, Auf den geweihten Tönen hat getragen, Wie muß jetzt fromme Kunst dein Schweigen klagen! Es ward Gemüthern, die von Erz, Durch dich die ungewohnte Sanftmuth theuer; Es war kein Ohr so stumpf, so roh kein Herz, Das nicht empfand, nicht glüht’ von heilgem Feuer; Und mächt’ger als sein Thron, ward Davids Leier!

It told the triumphs of our King, It wafted glory to our God; It made our gladden’d valleys ring, The cedars bow, the mountains nod; Its sound aspired to heaven and there abode! Since then, though heard on earth no more, Devotion and her daughter Love Still bid the bursting spirit soar To sounds that seem as from above, In dreams that day’s broad light can not remove.

Es feierte den Ewigen ihr Klang; Ihr Ton war seiner großen Werke Spiegel; Es horchten froh die Thäler dem Gesang, Die Zedern ließ er hüpfen und die Hügel, Und drang zu Gott auf der Begeistrung Flügel. Verklungen ist er jedem Ohr, Von Andacht nur und Liebe noch vernommen. Für diese bricht ihr hoher Geist hervor. In Tönen, Träumen, die vom Himmel kommen, Und die kein Tageslicht entführt dem Herz der Frommen.

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1139-2296 (2015-05-25 19:56:38)

Traduko de la Angla poemo “The harp the monarch minstrel swept” de Lord George Gordon Byron (∗1788-01-22 – †1824-04-18) en la Germanan de Franz Theremin (∗1780-03-19 – †1846-09-26).

Arg-1139-2297 (2015-05-25 23:28:20)

http://www.poezio.net/version?poem-id=1139&version-id=2296,2297